Find the equilibrium price of the following:
(1)
Qd = -20 + 3p
Qd = 220 - 5p

(2)
Qd = -45 + 8p
Qd = 125 - 2p
1
Expert's answer
2021-07-22T13:43:59-0400

Question 1:

-20 + 3p = 220 - 5p

5p + 3p = 220 + 20

8p = 240

p = 30

The equilibrium price is: 30


Question 2:

-45 + 8p = 125 -2p

2p + 8p = 125 + 45

10p = 170

p = 17

The equilibrium price is 17